# Limitations on Safe, Trusted, Artificial General Intelligence

### v1

@@ 3.2 Planning
In a planning problem we are given a sequence of states, a set of associated moves, a start state and a desired goal state. For any state $u$ and move pair, there is a an explicit program (which is provided as part of the problem specification) which returns the next state (certain moves may be illegal and may return in ‘not allowed’ states). The task is to find a sequence of moves which end up in the goal state from the start state, or to prove that it is not possible to reach the goal state from the start state.
@@
If a deterministic AI system is safe and trusted, then it cannot be an AGI system for planning. In particular, for such a system there is a planning problem instance for which the system outputs ‘don’t know’ but there is a short proof that the planning problem has no winning moves.
@@ 3.2.1 Halting for a specific program input instance
If a deterministic system is safe and trusted, then it cannot be an AGI system for the task of determining whether a program halts on a specific input instance.
We note that Theorem 3.6 follows from Theorem 3.8. This is because we can reduce the halting problem for a program-input pair to a planning instance. Given a program $P$ and input $I$, we construct a planning problem where the states correspond to the configurations of $P$ during its execution on $I$, and the moves represent single-step transitions between these configurations. The start state is the initial configuration of $P$ on input $I$, and the goal state is a halting configuration of $P$. The planning task is to determine whether a sequence of moves exists that leads from the start state to the goal state—this is equivalent to determining whether $P$ halts on $I$. Hence, if a safe and trusted system could solve all such planning instances, it would be able to solve the halting problem in Definition 3.7, contradicting Theorem 3.8. This establishes that, under our definitions, a safe and trusted system cannot be an AGI system for planning.
We now prove Theorem 3.8.
Proof of Theorem 3.8.
Algorithm 2 Turing_program
1:procedure $\mathtt{Turing\_program}$(Program $P$, Input II)
2:  if $A(P,I)$ == ‘does not halt’ then
3:   return 0
4:  else
5:   while $\mathtt{true}$ do $\triangleright$ run indefinitely
6:   end while
7:  end if
8:end procedure
We define
Now consider $\mathtt{self\_Turing\_program}(\mathtt{self\_Turing\_program})$.
Lemma 3.9.
If $A$ is safe then $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt, but the AI system $A$ cannot prove that it does not halt.
Proof.
Note that if $A$ is safe then the program cannot halt because of entering the if block in line 3, since the program only enters this block if the safe system $A$ determines that $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt. If $A$ does not determine that the program does not halt, then the program enters the infinite loop and never halts. Therefore, $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt if $A$ is safe.
The second part of the lemma has a similar proof. If $A$ determines that the program input pair $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt, then it does halt and we have a contradiction since $A$ is safe. Therefore if $A$ is safe then $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt, but $A$ cannot determine that the program input pair $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt.
Finally, note that the assumption of $A$ being safe is satisfied for a trusted system $A$, therefore for a trusted system we have a short proof that $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt, even though the system cannot solve this instance if it is safe. Hence the system cannot be an AGI system if it is safe and trusted. ∎
In Appendix A.1, we prove a similar version of Theorem 3.8 where the program provably halts on the input, but the AI system $A$ cannot determine so. This proof requires an additional assumption that $A$ is also well-behaved, i.e. it always terminates on an input, which can be ensured by having a fixed time limit on the execution of $A$. Note that since determining halting on a specific program input instance reduces to planning, this shows that for a safe, trusted and well-behaved system there are planning instances where humans can provably find a feasible plan, but the system will not be able to solve the instance.
@@ 3.3 Graph reachability
We show that safe, trusted AI systems need time almost as large as the size of the considered graph to solve certain reachability instances which actually admit a simple, solution. As in Section 3.2, we consider deterministic AI systems in this section for ease of exposition. In Appendix A.2, we extend to randomized AI systems.
@@
For any $T>0$, a fixed constant $c$, and any safe, trusted, deterministic AI system, there is a graph reachability problem instance of size $T$, for which the safe, trusted, deterministic AI system outputs ‘don’t know’ if it is run for time at most TcT-c, but there is a short, constant-sized proof that the answer is ‘not reachable’.
@@ 3.3.1 Time-bounded halting
If a deterministic system $A$ is safe and trusted, then it cannot be an AGI system for time-bounded halting. Specifically for a deterministic, safe, trusted system $A$ and for any $T>0$ and a fixed constant $c$, there is a program for which there is a short, constant-sized proof that it does not halt in $T$ steps, but $A$ will output ‘don’t know’ if it runs for time at most TcT-c.
@@
If $A$ is safe, then $\mathtt{self\_Turing\_T(self\_Turing\_T)}$ does not halt in time $T$. Moreover, for a fixed constant $c$, if $A$ is safe and is run for time at most TcT-c then $A$ will output ‘don’t know’ on whether $\mathtt{self\_Turing\_T(self\_Turing\_T)}$ halts in time at most $T$.
@@
If $\mathtt{self\_Turing\_T(self\_Turing\_T)}$ halts, it can only be because it enters the if block in line 3. However, it only enters this block if $A$ determines that it does not halt in time $T$. Since $A$ is safe, if the program enters the if block in line 3 then it must not halt in time $T$, and hence $\mathtt{self\_Turing\_T(self\_Turing\_T)}$ cannot halt in time $T$.
@@
Note that the execution of steps 2 and 3 of the program only take some fixed constant $c$ steps outside the execution of $A$ on $(\mathtt{self\_Turing\_T},\mathtt{self\_Turing\_T},T)$. Therefore if the AI system $A$ runs for time $T^{\prime}$ and outputs that $\mathtt{self\_Turing\_T(self\_Turing\_T)}$ does not halt in time $T$, then $\mathtt{self\_Turing\_T(self\_Turing\_T)}$ halts in time $T^{\prime}+c$. If $T^{\prime}<T-c$, then the program does halt in total time $T$, which contradicts safety. Therefore, a safe AI system $A$ must output ‘don’t know’ if it runs for time at most TcT-c, for some fixed constant $c$. ∎
@@
As in the previous proofs, note that the assumption of $A$ being safe is satisfied for a trusted system $A$, therefore for a trusted system we have a short proof that $\mathtt{self\_Turing\_T}$($\mathtt{self\_Turing\_T}$) does not halt in time $T$, even though $A$ cannot solve this instance if it is safe and run for time at most TcT-c. ∎
At the end of Section 3.2.1, we discussed an additional result about planning in the case where a feasible plan exists. We also show a similar result for graph reachability. In Appendix A.2, we prove a similar version of Theorem 3.13 under an additional assumption that $A$ always terminates in time $T$. We show that for such a system $A$ there is an instance which provably halts in time $T+c$ (for some constant $c$) if $A$ is safe, but the safe system $A$ cannot determine so. As before, since determining halting within a fixed time limit reduces to graph reachability on finite-sized graphs, this shows that for a safe, trusted system with an upper bound on its running time, there are graph reachability instances where humans can provably find a path, but the system will not be able to solve the instance in time slightly less than the size of the graph.
4 Impossibility result for calibration as the notion of safety
@@
In this section, we define a relaxed notion of safety, which we call calibration-safety. The notion is derived from the usual notion of calibration, a well-studied notion for ensuring reliability of a model’s prediction (Dawid, 1982; Van Calster et al., 2019).
Definition 4.1 (Calibration-Safety for Program Verification).
We define a system $A$ to be calibration-safe if for any program $P$ with input $I$:
@@
Note that similar to the definition of safety (Definition 1.2), calibration-safety does not put any requirement on the system if it decides to abstain with probability 1 on a given input. We show that if a system is calibration-safe, and in addition is also well-behaved, then it fails on certain instances which provably terminate with good probability.
@@
If the AI system $A$ is well-behaved and calibration-safe for program verification, then there is a program $P$ which provably halts with probability at least $0.99$, but $A$ abstains with probability 1 on the program $P$.
We note that Theorem 4.2 implies a similar impossibility result as Theorem 3.2 but with a relaxed notion of safety and a corresponding notion of trust. Similar to how we define trust in Definition 1.3 as the assumption that the system is safe, we can define calibration-trust as the assumption that the system is calibration-safe. Then Theorem 4.2 implies that for a well-behaved, calibration-safe and calibration-trusted system $A$, there is a program which can be proven to halt with probability at least 0.99, but $A$ will abstain with probability 1 on the program. Therefore, a well-behaved, calibration-safe and calibration-trusted system cannot be an AGI system.
@@ 5 Discussion
Our results show that safety, trust and AGI are mutually incompatible. We further discuss implications of the result and some possible critiques and clarifications.
@@
  • Worst-case nature of the results: While we demonstrate specific task instances which are not solvable by certain systems, we note that the system could still solve a vast number of interesting instances. However, the result still points to certain barriers which cannot be overcome by safe, trusted systems. Given the significant interest and economic capital being devoted to building safe AGI or superintelligent systems, we believe it is important to understand the barriers fundamental to any such technology. Somewhat more speculatively, note that our constructions rely on self-referential calls to the AI system, and when systems have general-purpose capabilities, such calls may not be implausible.
  • Limitations of human reasoning: We note that there is a long line of work on studying the limitations of human reasoning in cognitive science and other fields, and it has long been emphasized that human reasoning is resource-bounded and error-prone (Simon, 1957; Tversky and Kahneman, 1974). However, our goal is not to argue for strict superiority of human reasoning over AI, but to show a separation: for safe, trusted AI systems there are instances that humans can solve, but which are not solvable by the system.
@@ Acknowledgement
We thank Prabhakar Raghavan and Mukund Raghothaman for insightful discussions. VS was supported by an NSF CAREER Award CCF-2239265, an Amazon Research Award, a Google Research Scholar Award and an Okawa Foundation Research Grant. The work was done in part while VS was visiting the Simons Institute for the Theory of Computing. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any funding agencies such as the National Science Foundation, Amazon or Google.
@@ Appendix A Additional results
A.1 Impossibility of solving planning when a feasible plan exists
@@
If the AI system $A$ is safe and well-behaved for determining halting on a specific program input pair, then there is a program input instance pair for which there is a short proof that the instance always terminates, but $A$ cannot determine that the instance always terminates (the probability of $A$ giving the answer ‘always halts’ is 0).
Note that the condition ‘If the AI system $A$ is safe’ is satisfied when $A$ is trusted, and hence for a safe, trusted, well-behaved AI system there are program, input instances which the system cannot solve, but for which there is a short proof that the instance terminates. We now prove Theorem A.2.
@@
If $A$ is safe and well-behaved then $\mathtt{self\_Turing\_program\_v2(self\_Turing\_program\_v2)}$ always halts, but the safe AI system $A$ cannot determine that it always halts.
@@
Note that the if condition in step 2 always terminates if $A$ is well-behaved. Suppose $A$ outputs ‘always halts’ on some randomness. Then, $\mathtt{self\_Turing\_program\_v2(self\_Turing\_program\_v2)}$ does not halt on some randomness. If $A$ is safe, then this is a contradiction. Therefore, if $A$ is safe then it must output ‘always halts’ with 0 probability.
Note that if $A$ does not output that $\mathtt{self\_Turing\_program\_v2(self\_Turing\_program\_v2)}$ ‘always halts’, then the program enters the else condition and immediately terminates, and therefore halts. Therefore if $A$ is safe and well-behaved, then $\mathtt{self\_Turing\_program\_v2(self\_Turing\_program\_v2)}$ always halts.
A.2 Impossibility of solving feasibility when a path exists in the graph
We prove a similar result to Theorem 3.13 in this section, for the case where the program terminates on the given input within some time bound. As in Appendix A.1, we also strengthen the result to allow randomized AI systems. We first extend Definition 3.12 to allow for randomized programs.

### v2

@@ 3.2 Planning
In a planning problem we are given a sequence of states, a set of associated moves, a start state and a desired goal state. For any state and move pair, there is an explicit program (which is provided as part of the problem specification) which returns the next state (certain moves may be illegal and may return in ‘not allowed’ states). The task is to find a sequence of moves which end up in the goal state from the start state, or to prove that it is not possible to reach the goal state from the start state.
@@
If a deterministic AI system is accurate and trusted, then it cannot be a human-level reasoning system for planning. In particular, for such a system there is a planning problem instance for which the system outputs ‘don’t know’ but there is a short proof that the planning problem has no winning moves.
@@ 3.2.1 Halting for a specific program input instance
If a deterministic system is accurate and trusted, then it cannot be a human-level reasoning system for the task of determining whether a program halts on a specific input instance.
We note that Theorem 3.6 follows from Theorem 3.8. This is because we can reduce the halting problem for a program-input pair to a planning instance. Given a program $P$ and input $I$, we construct a planning problem where the states correspond to the configurations of $P$ during its execution on $I$, and the moves represent single-step transitions between these configurations. The start state is the initial configuration of $P$ on input $I$, and the goal state is a halting configuration of $P$. The planning task is to determine whether a sequence of moves exists that leads from the start state to the goal state—this is equivalent to determining whether $P$ halts on $I$. Hence, if an accurate and trusted system could solve all such planning instances, it would be able to solve the halting problem in Definition 3.7, contradicting Theorem 3.8. This establishes that, under our definitions, an accurate and trusted system cannot be a human-level reasoning system for planning.
The proof of Theorem 3.8 appears in Appendix A.1, and is similar to the proof of Theorem 3.2, and also the next result, Theorem 3.12. In Appendix A.2, we prove a similar version of Theorem 3.8 where the program provably halts on the input, but the AI system $A$ cannot determine so. This proof requires an additional assumption that $A$ is also well-behaved, i.e. it always terminates on an input, which can be ensured by having a fixed time limit on the execution of $A$. Note that since determining halting on a specific program input instance reduces to planning, this shows that for an accurate, trusted and well-behaved system there are planning instances where humans can provably find a feasible plan, but the system will not be able to solve the instance.
@@ 3.3 Graph reachability
We show that accurate, trusted AI systems need time almost as large as the size of the considered graph to solve certain reachability instances which actually admit a simple solution. As in Section 3.2, we consider deterministic AI systems in this section for ease of exposition. In Appendix A.3, we extend to randomized systems.
@@
For any $T>0$, a fixed constant $c$, and any accurate, trusted, deterministic AI system, there is a graph reachability problem instance of size $T$, for which the accurate, trusted, deterministic AI system outputs ‘don’t know’ if it is run for time at most TcT-c, but there is a short, constant-sized proof that the answer is ‘not reachable’.
@@ 3.3.1 Time-bounded halting
If a deterministic system $A$ is accurate and trusted, then it cannot be a human-level reasoning system for time-bounded halting. Specifically for a deterministic, accurate, trusted system $A$ and for any $T>0$ and a fixed constant $c$, there is a program for which there is a short, constant-sized proof that it does not halt in $T$ steps, but $A$ will output ‘don’t know’ if it runs for time at most TcT-c.
@@
If $A$ is accurate, then $\mathtt{self\_Turing\_T(self\_Turing\_T)}$ does not halt in time $T$. Moreover, for a fixed constant $c$, if $A$ is accurate and is run for time at most TcT-c then $A$ will output ‘don’t know’ on whether $\mathtt{self\_Turing\_T(self\_Turing\_T)}$ halts in time at most $T$.
@@
If $\mathtt{self\_Turing\_T(self\_Turing\_T)}$ halts, it can only be because it enters the if block in line 3. However, it only enters this block if $A$ determines that it does not halt in time $T$. Since $A$ is accurate, if the program enters the if block in line 3 then it must not halt in time $T$, and hence $\mathtt{self\_Turing\_T(self\_Turing\_T)}$ cannot halt in time $T$.
@@
Note that the execution of steps 2 and 3 of the program only take some fixed constant $c$ steps outside the execution of $A$ on $(\mathtt{self\_Turing\_T},\mathtt{self\_Turing\_T},T)$. Therefore if the AI system $A$ runs for time $T^{\prime}$ and outputs that $\mathtt{self\_Turing\_T(self\_Turing\_T)}$ does not halt in time $T$, then $\mathtt{self\_Turing\_T(self\_Turing\_T)}$ halts in time $T^{\prime}+c$. If $T^{\prime}<T-c$, then the program does halt in total time $T$, which contradicts accuracy. Therefore, an accurate AI system $A$ must output ‘don’t know’ if it runs for time at most TcT-c, for some fixed constant $c$. ∎
@@
As in the previous proof, if $A$ is accurate and trusted by $H$, then individuals $h\in H$ assume that $A$ is accurate and can hence prove that $\mathtt{self\_Turing\_T}$($\mathtt{self\_Turing\_T}$) does not halt in time $T$, even though $A$ cannot solve this instance if it is accurate and run for time at most TcT-c. ∎
At the end of Section 3.2.1, we discussed an additional result about planning in the case where a feasible plan exists. We also show a similar result for graph reachability, discussed in Appendix A.3.
4 Impossibility Result for Other Semantic Properties
In this section, we extend the program verification result from halting to other properties of programs. This includes properties such as verifying if a program executes certain pre-defined “harmful” behavior, which could be important from the perspective of safety. More formally, we first define the task of determining if an input program executes certain harmful behavior.
Definition 4.1 (Harmful state execution).
A program is “harmless” if it does not enter certain pre-defined “harmful” states during execution, and “harmful” otherwise.
We show an impossibility result for this property which is analogous to our result for verifying if an input program halts.
Theorem 4.2.
Consider an AI system $A$ which is well-behaved and itself does not enter pre-defined “harmful” states during its execution. Then if $A$ is accurate and trusted for verifying harmful state execution (Definition 4.1), then it cannot be a human-level reasoning system.
This result follows as a corollary of a more general theorem, which extends our result for halting to any non-trivial, semantic property of a program. These are the same conditions under which Rice’s theorem extends Turing’s undecidability result (Rice, 1953).
Definition 4.3 (Non-trivial, semantic property (Rice, 1953)).
A semantic property is a property of program which concerns its behavior (e.g. “does the program always terminate?”) as opposed to its syntax (e.g. “does the program have a while loop?”). A non-trivial property is a property which is neither true for all programs, nor false for all programs.
Examples of semantic properties that we have already discussed are halting and harmful state execution. Another semantic property is correctness, for example if a program intended to determine if an input number is prime correctly outputs whether the number is prime.
We now state the result, proved in Appendix A.4. The result applies to AI systems $A$ which are well-behaved, and whose execution does not trivially lead to the desired property $\pi$ being satisfied. Since our constructions are self-referential, this condition ensures that the program does not automatically have the property $\pi$ by virtue of the execution of $A$. For instance, if $A$ itself always entered “harmful” states during its execution, then a program which always calls $A$ also trivially enters “harmful” states. Note that from the perspective of safety, if $A$ itself entered “harmful” states then it would also have unsafe behavior.
Theorem 4.4.
Consider any non-trivial, semantic property $\pi$ of programs. Consider an AI system $A$ which is well-behaved and has the property that if some program $P$ calls $A$ during its execution then the execution of $A$ never automatically leads to the property $\pi$ being satisfied for program $P$. Then if $A$ is accurate and trusted for the task of verifying if an input program has property $\pi$, then it cannot be a human-level reasoning system for this task.
5 Impossibility Result for Calibration
@@
In this section, we define a relaxed notion of accuracy. The notion is derived from the usual notion of calibration, a well-studied notion for ensuring reliability of a model (Dawid, 1982; Van Calster et al., 2019).
Definition 5.1 (Calibration for Program Verification).
We define a system $A$ to be calibrated for program verification if for any program $P$ with input $I$:
@@
Note that similar to the definition of accuracy (Definition 1.2), calibration does not put any requirement on the system if it decides to abstain with probability 1 on a given input. We show that if a system is calibrated, and in addition is also well-behaved, then it fails on certain instances which provably terminate with good probability.
@@
If the AI system $A$ is well-behaved and calibrated for program verification, then there is a program $P$ which provably halts with probability at least $0.99$, but $A$ abstains with probability 1 on $P$.
We note that Theorem 5.2 implies a similar impossibility result as Theorem 3.2 but with a relaxed notion of accuracy and a corresponding notion of trust. In the context of calibration, trust in Definition 1.3 is the assumption that the system is calibrated. Then Theorem 5.2 implies that for a well-behaved, calibrated and trusted system $A$, there is a program which can be proven to halt with probability at least 0.99, but $A$ will abstain with probability 1 on the program. Therefore, a well-behaved, calibrated and trusted system cannot be a human-level reasoning system.
@@ 6 Discussion
Our results show that accuracy, trust and human-level reasoning are mutually incompatible. We further discuss implications of the result and some possible critiques and clarifications.
  • Worst-case nature of the results: While we demonstrate specific task instances which are not solvable by certain systems, we note that the system could still solve a vast number of interesting instances. However, the result still points to certain barriers which cannot be overcome by accurate, trusted systems. Given the significant interest and economic capital being devoted to building accurate or reliable human-level reasoning, we believe it is important to understand the barriers fundamental to any such technology. By way of analogy, Gödel’s and Turing’s results pointed to fundamental barriers to mathematics and computation. Though these barriers were worst-case, they identified the limits of what is possible and what is not possible. Similarly, we believe that it is important to outline the limits of what is possible in the context of AI and requirements of accuracy and trust. Somewhat more speculatively, note that our constructions rely on self-referential calls to the AI system, and when systems have general-purpose capabilities, such calls may not be implausible.
@@
  • Limitations of human reasoning: We note that there is a long line of work on studying the limitations of human reasoning in cognitive science and other fields, and it has long been emphasized that human reasoning is resource-bounded and error-prone (Simon, 1957; Tversky and Kahneman, 1974). However, our goal is not to argue for strict superiority of human reasoning over AI, but to show a separation: for accurate, trusted AI systems there are instances that humans can solve, but which are not solvable by the system.
@@ Acknowledgement
We thank Prabhakar Raghavan, Mukund Raghothaman and anonymous reviewers for insightful discussions and helpful feedback. VS was supported by NSF awards 2239265, 2346058 and an Okawa Foundation Research Grant. The work was done in part while VS was visiting the Simons Institute for the Theory of Computing. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any funding agencies such as the National Science Foundation.
@@ Appendix A Additional results
A.1 Proof of Theorem 3.8
See 3.8
Proof of Theorem 3.8.
Algorithm 3 Turing_program
1:procedure $\mathtt{Turing\_program}$(Program $P$, Input II)
2:  if $A(P,I)$ == ‘does not halt’ then
3:   return 0
4:  else
5:   while $\mathtt{true}$ do $\triangleright$ run indefinitely
6:   end while
7:  end if
8:end procedure
Now define
We consider:
$\mathtt{self\_Turing\_program}(\mathtt{self\_Turing\_program})$
Lemma A.1.
If $A$ is accurate then $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt, but the AI system $A$ cannot prove that it does not halt.
Proof.
Note that if $A$ is accurate then the program cannot halt because of entering the if block in line 3, since the program only enters this block if the accurate system $A$ determines that $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt. If $A$ does not determine that the program does not halt, then the program enters the infinite loop and never halts. Therefore, $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt if $A$ is accurate.
The second part of the lemma has a similar proof. If $A$ determines that the program input pair $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt, then it does halt and we have a contradiction since $A$ is accurate. Therefore if $A$ is accurate then $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt, but $A$ cannot determine that the program input pair $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt.
Finally, if $A$ is accurate and trusted by $H$, then individuals $h\in H$ assume that $A$ is accurate and can hence prove that $\mathtt{self\_Turing\_program(self\_Turing\_program)}$ does not halt, even though the system cannot solve this instance if it is accurate. This completes the proof of the theorem.
A.2 Impossibility of solving planning when a feasible plan exists
@@
If the AI system $A$ is accurate and well-behaved for determining halting on a specific program input pair, then there is a program input instance pair for which there is a short proof that the instance always terminates, but $A$ cannot determine that the instance always terminates (the probability of $A$ giving the answer ‘always halts’ is 0).
Note that if $A$ is accurate and trusted by $H$, then individuals $h\in H$ assume that $A$ is accurate and hence for an accurate, trusted, well-behaved AI system there are program, input instances which the system cannot solve, but for which there is a short proof (i.e., the proof of Theorem A.3 since the condition of $A$ being accurate is satisfied under trust) that the instance terminates. We now prove Theorem A.3.
@@
If $A$ is accurate and well-behaved then $\mathtt{self\_Turing\_program\_v2(self\_Turing\_program\_v2)}$ always halts, but the accurate AI system $A$ cannot determine that it always halts.
@@
Note that the if condition in step 2 always terminates if $A$ is well-behaved. Suppose $A$ outputs ‘always halts’ on some randomness. Then, $\mathtt{self\_Turing\_program\_v2(self\_Turing\_program\_v2)}$ does not halt on some randomness. If $A$ is accurate, then this is a contradiction. Therefore, if $A$ is accurate then it must output ‘always halts’ with 0 probability.
Note that if $A$ does not output that $\mathtt{self\_Turing\_program\_v2(self\_Turing\_program\_v2)}$ ‘always halts’, then the program enters the else condition and immediately terminates, and therefore halts. Therefore if $A$ is accurate and well-behaved, then $\mathtt{self\_Turing\_program\_v2(self\_Turing\_program\_v2)}$ always halts.
A.3 Impossibility of solving feasibility when a path exists in the graph
We prove a similar result to Theorem 3.12 in this section, for the case where the program terminates on the given input within some time bound. The result is shown under an additional assumption that $A$ always terminates in time $T$. We show that for such a system $A$ there is an instance which provably halts in time $T+c$ (for some constant $c$) if $A$ is accurate, but the accurate system $A$ cannot determine so. As before, since determining halting within a fixed time limit reduces to graph reachability on finite-sized graphs, this shows that for an accurate, trusted system with an upper bound on its running time, there are graph reachability instances where humans can provably find a path, but the system will not be able to solve the instance in time slightly less than the size of the graph.
As in Appendix A.2, we also strengthen the result to allow randomized AI systems. We first extend Definition 3.11 to allow for randomized programs.